OpenStudy (anonymous):

which of the following could be points on the unit circle A)( root3/2 , 1/3) B) (-2/3 , root 5/30 C) (0.8, -0.6) D) (1,1)

OpenStudy (math&ing001):

The unit circle has a radius of r=1, so check for pairs that verify \(x ^{2}+y ^{2}=1\).

OpenStudy (anonymous):

0.8^2 + 0.6^2 = 0.64 + 0.36 =1 So C